OpenStudy (anonymous):

Can someone please help me with some Special Right Triangles? I will post them in a link below.

OpenStudy (anonymous):

http://gyazo.com/89545e2b6570d8819ed399807c220b26 The first one.

OpenStudy (luigi0210):

Do you know the special right triangles?

OpenStudy (anonymous):

I don't, sorry..

OpenStudy (anonymous):

@whpalmer4 Could you maybe help me again?

OpenStudy (luigi0210):

On a 45-45-90 triangle, both the legs have the same value, while the hypotenuse has same value but with a square root of 2 value.. coming from \[\LARGE a^2+b^2=c^2\]

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

So...??

OpenStudy (luigi0210):

whpalmer is here o:

OpenStudy (whpalmer4):

So plug in the values and solve it :-) You know that one side is 8. Let that side be \(a\). You know that the other side is equal, because you have an isosceles triangle (2 angles are congruent). Let that side be \(b\) which also equals 8. That gives you: \[a^2+b^2=c^2\]\[8^2+8^2=c^2\] Can you solve that for \(c\)?

OpenStudy (anonymous):

So would it be 16?

OpenStudy (whpalmer4):

No... \[8^2 = 8*8 = \]

OpenStudy (anonymous):

Uh..I got a big number from that but it's not there on my answers list?

OpenStudy (whpalmer4):

What is 8*8? Don't jump ahead to thinking you have the answer!

OpenStudy (anonymous):

64. So it would be 128.

OpenStudy (whpalmer4):

\[8^2 + 8^2 = c^2\]\[128 = c^2\]What does \(c=\)

OpenStudy (whpalmer4):

Put another way:\[c*c = 128\]\[c = \]

OpenStudy (anonymous):

128 right? I'm so confused, sorry.

OpenStudy (whpalmer4):

128*128 = 128?

OpenStudy (luigi0210):

And if I'm allowed to jump it.. \[\LARGE \sqrt{64*2}=c\]c:

OpenStudy (anonymous):

Oh! I thought I did that, oops.

OpenStudy (whpalmer4):

Yes, @Luigi0210 you are always welcome to jump in!

OpenStudy (anonymous):

16384

OpenStudy (whpalmer4):

yes, 128*128 = 16384. We want a number,\(c\), such that \(c*c = 128\). You've shown that \(c\) is definitely not 128 :-)

OpenStudy (anonymous):

@Luigi0210 Thank you for your help too!!!

OpenStudy (anonymous):

Okay..

OpenStudy (whpalmer4):

If \[c^2 = 128\]we take the square root of both sides: \[\sqrt{c^2} = \sqrt{128}\]What is the square root of \(c^2\), assuming that \(c>0\)?

OpenStudy (whpalmer4):

Hint: \[\sqrt{c^2} = \sqrt{c*c}\]

OpenStudy (anonymous):

0?

OpenStudy (whpalmer4):

perhaps you don't understand the concept of a square root? The square root of some number \(x\) is another number, \(a\), such that \(a*a = x\). The square root of 4 is 2, because 2*2 = 4. The square root of 9 is 3, because 3*3 = 9. Can you tell me what the square root of 64 is?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

I understand the concept, it just looks like soup to me.

OpenStudy (whpalmer4):

Sigh. 4*4 = 5*5 = 6*6 = 7*7 = 8*8 = 9*9 = 10*10 = What number times itself gives you 64? The answer is in the table you just made.

OpenStudy (anonymous):

(Sorry if it looks like I'm gone, bad internet connection.)

OpenStudy (anonymous):

I just said 8.

OpenStudy (anonymous):

Sorry.

OpenStudy (whpalmer4):

Okay. So \[\sqrt{64} = 8\]However, we need to find the square root of 128, or\[\sqrt{128} = \sqrt{64*2}\]as @Luigi0210 helpfully pointed out. By the properties of square roots, we can write \[\sqrt{128} = \sqrt{64*2} = \sqrt{64}*\sqrt{2}\]We know that \(\sqrt{64} = 8\), so that becomes \[\sqrt{128} = \sqrt{64*2} = \sqrt{64}*\sqrt{2} = 8*\sqrt{2}\] Is there an integer that we can square to get 2? If not, we're done.

OpenStudy (anonymous):

Unless four would work, no?

OpenStudy (whpalmer4):

Does 4*4 = 2? You're confusing squaring with taking a square root. No, there is no rational number we can square (multiply by itself) to get 2. \[1*1=1\]\[2*2=4\]If such a number existed, it would be somewhere between 1 and 2. Our final answer is \[c =\sqrt{128} = \sqrt{64}*\sqrt{2} = 8*\sqrt{2} = 8\sqrt{2}\] What @Luigi0210 was trying to tell you with the reference to special triangles is that in an isosceles right triangle, the ratio of the length of the hypotenuse to the length of either side is always \(\sqrt{2}:1\). So if you know that the side is \(8\), you can immediately conclude that the hypotenuse is \(8\sqrt{2}\). If the side is 10, the hypotenuse is \(10\sqrt{2}\). It also works the other way, of course: if you know what the hypotenuse is, you can find the side by dividing by \(\sqrt{2}\). This particular relationship is ONLY true if you have a 45/45/90 triangle. I would encourage you to watch this video: https://www.khanacademy.org/math/arithmetic/exponents-radicals/radical-radicals/v/understanding-square-roots You seem to be on rather unsteady ground when it comes to square roots. Maybe the video can help clear it up. If not, ask your teacher for a review.